# thanks: equal contributionthanks: equal contribution

### v1

@@ IV Filamentous microbes
Figure 4: Examples of cyanobacteria self-organisation. (a) Many cyanobacteria species grow into filaments made out of a chain of identical cells Culture Collection of Algae, SAG . (b) These filaments are capable of organizing into intricate structures like reticulate mats Faluweki and Goehring (2022), spirals Gong and Prakash (2023) (c), and even ropes Garcia-Pichel and Wojciechowski (2009) (d).
@@
Various microbial organisms have evolved a filamentous body shape. Examples of such organisms that have attracted the interest of physicists include various filamentous cyanobacteria Faluweki et al. (2023); Faluweki and Goehring (2022); Kurjahn et al. (2024b, a); Cammann et al. (2024); Ehlers et al. (1996), filamentous phages Dogic and Fraden (1997, 2006); Purdy and Fraden (2004), and nematodes, Bilbao et al. (2013); Gart et al. (2011); Backholm et al. (2013); Niebur and Erdös (1991) that straddle the boundary between the microscopic and macroscopic world. As there are too many filamentous microorganisms to list, this section will focus on discussing more generally the implications of a filamentous body shape at the microscale for a few important classes of microorganisms.
@@
Having a body with a large aspect-ratio like a filament comes with some advantages over microbes shaped as spheres or short rods. An elongated shape increases the chance of meeting one another in disordered environments Arguedas-Leiva et al. (2022). Thanks to a filamentous shape, microbial organisms can self-propel without appendages. Rather than using flagella or cilia to generate motility, they can bend their whole body to generate thrust thanks to the different drag coefficients that a slender body experiences in the direction tangential and perpendicular to its motion in a viscous fluid Lauga and Powers (2009); Hancock (1953) (see Box III). A filamentous shape has been observed to be a viable strategy to avoid predation by protists by simply being too long to be ingested easily Pernthaler (2005) (this approach is however futile at a certain size difference between devourer and the devoured, as is evident in the popularity of spaghetti as a human food source). The increased surface-to-volume ratio of a filamentous shape allows for easier uptake of nutrients Young (2006), but also allows stronger attachment to substrates Weiss et al. (1995) in bacteria, because of a larger contact patch, compared to a spherical one.
@@
A filament’s most crucial mechanical property is its bending stiffness. It is known one can draw conclusions about the forces at play from purely geometrical observations (see Box II). Techniques to measure the bending stiffness $\beta_{\mathrm{f}}$ and Young’s modulus $E$ of gliding filaments with the help of microfluidics Faluweki and Goehring (2022) and mircopipette techniques Kurjahn et al. (2024a) have been developed (see Table 1).
@@
Filamentous cyanobacteria often exist in a surface-bound but motile state, called gliding Hoiczyk (2000); Wilde and Mullineaux (2015); Tchoufag et al. (2019). The underlying mechanism that allows the generation of propulsive forces in this state is still a matter of debate McBride (2001); Read et al. (2007); Dhahri et al. (2013); Koiller et al. (2010); Halfen and Castenholz (1970), as no cyanobacteria species are known to possess flagella Wilde and Mullineaux (2015). The main mechanisms that are debated are: type IV pili extension and retraction Khayatan et al. (2015); Wilde and Mullineaux (2015), fibrils Halfen and Castenholz (1970); Read et al. (2007), ejection of a jet of extracellular slime from junctional pores Hoiczyk and Baumeister (1995, 1998); Wolgemuth and Oster (2004); Dhahri et al. (2013) (at least for some species), and acoustic streaming Ehlers et al. (1996); Koiller et al. (2010); Ehlers and Koiller (2011). By observing the self-buckling of a filament, the forces at play in its propulsion can be inferred via the bending stiffness Kurjahn et al. (2024a). Such observations combined with the independence of propulsion speed on filament length implies, that all cells within a cyanobacteria filament contribute roughly equal to its propulsion.
@@
Cyanobacteria locomotion exhibits another interesting feature: path-tracking motion, where the body follows the path laid by the forward end of the filament Faluweki et al. (2023). This type of locomotion in worms is also known as metameric locomotion Du et al. (2022). Physically, information about a filament’s orientation persists over the time it takes a filament to traverse its own length Du et al. (2022). Although their motility is induced by polar forces, filaments themselves are seemingly nematic, with no head or tail end. In fact, cyanobacteria have been observed to reverse their direction of motion frequently, as a strategy to navigate towards more favorable conditions, and even to form communities. Wilde and Mullineaux (2015); Rosko et al. (2024); Abbaspour et al. (2023); Pfreundt et al. (2023). Given the delocalised nature of the filaments propulsive mechanism, a high degree of intercellular coordination is required to reverse the direction of motion, with the mechanisms of communication between cells not yet understood. Confined active filaments with low bending rigidity have been shown to curl up into tightly wound spirals (see Fig. 4(c)), effectively trapping themselves, greatly impeding their mobility Duman et al. (2018); Prathyusha et al. (2018); Isele-Holder et al. (2015); Lin et al. (2014). Utilising reversals is a strategy to escape this trapped state Gong and Prakash (2023).
@@
Having a high aspect-ratio allows microbes to effectively sample changing conditions within space. While organisms with short bodies commonly have to move around to find favourable conditions, essentially sampling the environment at random to inform their movement through chemo- or phototaxis, elongated filaments can sample gradients along their body and react before they have wholly immersed themselves in an unfavourable environment. Cyanobacteria, for instance, have been observed to react to light shining only onto parts of their body Nultsch and Wenderoth (1983); Wilde and Mullineaux (2017), showing scotophobic responses, well before the whole filament loses access to light by reversing its gliding direction Kurjahn et al. (2024b).
@@
Having an elongated body may even allow organisms to be in two places with different conditions at once. Cable bacteria are filamentous, multicellular bacteria belonging to the family Desulfobulbaceae. They can conduct electrons over centimetre scales Pfeffer et al. (2012); Marzocchi et al. (2014). Through their filamentous shape they can couple sulfide oxidation in deeper, anoxic marine sediment with oxygen reduction in surface layers Nielsen et al. (2010); Bjerg et al. (2023); Risgaard-Petersen et al. (2015). Electrical activity is also involved in a microbial feat of engineering: some electroactive bacteria produce nanowires to transfer electrons to extracellular electron acceptors Reguera et al. (2005). These are, e.g.,the Geobacter sulfurreducens or the Shewanella oneidensis MR-1 bacteria El-Naggar et al. (2010); Pirbadian et al. (2014); Schkolnik et al. (2015); Logan et al. (2019).
@@
Filamentous microbes can explore spaces inaccessible to plumper organisms with the same volume. Semiflexible active filaments have been shown to effectively explore porous environments Mokhtari and Zippelius (2019). Many soil dwelling microbes are filamentous in shape, in addition to their ability to explore a porous environment it has been speculated that their shape may enable them to fix their position in space, by wrapping around soil particles Young (2006), allowing them to resist being washed away by percolating water. Entangled filamentous microbes have been observed to stabilize their habitats and protect them against erosion Kurtz Jr and Netoff (2001); Pluis (1994). Furthermore, nonmotile cells exploring a polymeric fluid can form long cables as they proliferate Corte et al. (2025).
@@
Their filamentous nature allows filamentous organisms such as cyanobacteria to organise into structures much larger than an individual filament Faluweki et al. (2023); Cammann et al. (2024); Shepard and Sumner (2010); Tamulonis and Kaandorp (2014). Their filamentous shape allows them to glide over one another without losing contact with the substrate. The dynamics of gliding filaments result in the formation of a network-like reticulate pattern of filament bundles Faluweki et al. (2023); Cammann et al. (2024); Tamulonis and Kaandorp (2014) (see Fig. 4(b)). Under different conditions, filaments have been observed to form aster-like aggregates Pfreundt et al. (2023). These relatively simple, mostly two-dimensional structures may serve as a starting point for filaments to organize themselves into even more intricate three-dimensional structures, ranging from complex patterned biomats Shepard and Sumner (2010); Cuadrado and Pan (2018) to the filaments weaving themselves into ropes Garcia-Pichel and Wojciechowski (2009) (see Fig. 4(d)), bridging large distances. The physical mechanisms behind this level of self-organisation and how the mechanical properties of these structures compare to that of a single filaments is yet to be understood, but can shed light on biofilm formation Mazza (2016); Sauer et al. (2022). It is plausible to presume that such complex assemblys could exhibit unusual rheological properties, just like those observed in cytosceletal networks (see Sec. II).
@@ V Undulatory locomotion at the macroscopic scale
At the organismal scale, animals with a slender and deformable body shape are also ‘active filaments’, and in some cases the analysis and framework presented in the previous sections could still apply. Diverse animals locomote by propagating undulatory waves. These body deformations are used to propel the organisms forward, although the actuation mechanisms may be very different across species and spatial scales (Figure 5). Well-studied examples include crawling worms Costa et al. (2024); Stephens et al. (2008), swimming eels Gillis (1998), and gliding snakes Jayne (2020). The study of both limbed and limbless movement has paved the way for understanding the origins and principles of neuro-mechanical control in diverse organisms, particularly the interactions between neural patterning and muscular activity Grillner et al. (1991). Undulatory or sinusoidal locomotion strategies evolved across different phyla, often according to a basic neural circuit design involving mutually inhibitory halves with the left and right sides deforming alternately Katz (2016). These insights into the neural basis of behaviour have led to the development of novel bio-inspired mechanisms for actuating synthetic systems including robots Ramdya and Ijspeert (2023).
@@
Since wave-like deformations are inherently non-reciprocal, they form the basis of diverse motility strategies through different media. In some cases, despite the large difference in scale compared to microscopic organisms such as flagellates and cyanobacteria (see Sec. III and IV), the undulatory locomotion of macroscopic animals such as nematodes and snakes can still be considered to operate in an overdamped regime where inertia is negligible, and where friction (instead of fluid viscosity) dominates. This is particularly true of locomotion in sand or granular media, where a resistive force theory based on drag anisotropy (see Box III) has been derived empirically, inspired by the movement of slender bodies through viscous fluids Hatton and Choset (2011); Zhang and Goldman (2014).
@@
Figure 5: Undulatory locomotion of active filaments across scales. (a) For C. elegans crawling movement on agar Padmanabhan et al. (2012). (b) Sidewinding locomotion of a snake, and (c) the timetrace of tracked points along the body of the snake as it moves on sand Rieser et al. (2021). (d) Vinegar eels self-organise into metachronal waves Peshkov et al. (2022). (e) Topological entanglement in worm blobs Patil et al. (2023). (f) Bio-inspired snake-like robots with CPG-like control modules Wu and Ma (2010).
@@
At the individual organism scale, distinct undulatory behaviours can arise from the same underlying structure and neural circuitry. A given body shape can be decomposed into a small number of principal components or oscillation modes, reflecting the intrinsically low-dimensional nature of the posture space, and allowing diverse undulatory shape modes to be reconstructed with only a small number of parameters Yemini et al. (2013). In the example of C. elegans crawling, its motion in 2D is stereotyped and well-captured by just six basis modes, or eigenworms Stephens et al. (2008); Costa et al. (2024). Frequency and amplitude thus uniquely define a gait, which in turn determines the distance traversed by the organism in one cycle. Notions such as geometric phase can help understand and predict the mapping between these gaits and eventual displacement in the lab frame Rieser et al. (2024). These dynamics are not totally unconstrained however, undulatory locomotion across swimming sperm, nematodes, and even fish in overdamped environments may be subject to a universal dispersion relationship Pierce et al. (2024); Sánchez-Rodríguez et al. (2023). Thus, active filaments can achieve highly reconfigurable dynamics with minimal neuromechanical coupling.
@@
Collectively, filamentous organisms also engage in novel behaviours that may not be accessible by individuals. At sufficiently high densities, these filaments self-assemble to form adaptive active materials with both solid-like and liquid-like properties, that are capable of emergent functions such as directed locomotion or smart navigation. Macroscopic ‘blobs’ of the 1 cm-long California blackworm Lumbriculus variegatus undergo dynamic collective movements to avoid desiccation or thermal stress without the need for any centralised control Ozkan-Aydin et al. (2021). The entanglement, or ‘braiding’, process enables fast and reversible tangling and untangling of worm collectives. Intriguingly, L. variegatus blobs can take minutes to assemble but only milliseconds to disentangle. A recent study proposed an explanation for this phenomenon by experimentally imaging the 3D trajectories of the worms and mapping the dynamics onto a 3D filament model of Kirchhoff rods Patil et al. (2023). The unique, activity-dependent rheology of entangled blobs of Tubifex tubifex worms was also studied and shown to be shear-thinning, but with a different scaling compared to polymer melts Deblais et al. (2020). These properties can even be harnessed to enable activity-dependent separation (chromatography) of worm mixtures Heeremans et al. (2022). In another free-swimming nematode Turbatrix aceti (vingar eels), concentrated suspensions of these organisms display large-scale metachronal waves surprisingly similar to those found in ensembles of cilia and flagella Peshkov et al. (2022); Quillen et al. (2021). Understanding how these deformable filamentous bodies actively control the topology and rheology of the collective could inspire novel designs of functionally tunable synthetic materials Deblais et al. (2023).
@@
More generally, undulatory or anguilliform locomotion in elongated bodies allows for direct interactions between an organism and the surrounding medium, which can yield novel principles of biomimetic/bioinspired locomotion reliant on embodied design with few mechanical components. Self-propulsion can also be easily adapted for movement through different environments, for example for snakes the same basic locomotion mechanism can be equally effective for locomotion on land as well as in water Taylor (1952); Stin et al. (2024). In more extreme cases, ‘flying’ snakes from the genus Chrysopelea use body undulations to generate lift to glide stably over several metres through the air Yeaton et al. (2020).
@@
From a control theoretic perspective, undulatory locomotion is universal perhaps because it reduces the need for more complex neural control mechanisms and can be achieved simply by coordinating alternating waves of (usually muscle) contraction and relaxation. In some cases, oscillations emerge spontaneously as a result of instabilities, for example dynamics reminiscent of ciliary beating have been observed in macroscopic chains of self-propelled microrobots, with the transition mapped to a supercritical Hopf bifurcation Zheng et al. (2023). Indeed substrate-attached, microscopic elastic filaments forced at one end also exhibit oscillations in different regimes De Canio et al. (2017); Bayly and Dutcher (2016); Man et al. (2020). The propagation of waves thus provides one of the simplest possible appendage or body actuation mechanisms, and naturally breaks time-reversal symmetry (particularly relevant for inertia-free swimming). This results in highly adaptable and flexible control of movement where both the speed and direction of travel can be readily changed in response to stimuli.

### v2

@@ IV Filamentous microbes
Figure 4: Examples of filamentous cyanobacteria self-organisation. (a) Many cyanobacteria species grow into filaments made out of a chain of identical cells [306]. These filaments are capable of organizing into intricate structures like (b) reticulate mats [87], (c) spirals [307], and (d) even ropes [308].
@@
Various microbial organisms have evolved a filamentous body shape. Examples of such organisms that have attracted the interest of physicists include various species of filamentous cyanobacteria [309, 87, 214, 89, 310, 311], filamentous phages [312, 313, 314], and nematodes, [315, 316, 90, 317] that straddle the boundary between the microscopic and macroscopic world. As there are too many filamentous microorganisms to list, this section will focus on discussing more generally the implications of a filamentous body shape at the microscale for a few important classes of microorganisms.
@@
Having a body with a large aspect-ratio like a filament comes with some advantages over microbes shaped as spheres or short rods. An elongated shape increases the chance of meeting one another in disordered environments [318]. Thanks to a filamentous shape, microbial organisms can self-propel without appendages. Rather than using flagella or cilia to generate motility, they can bend their whole body to generate thrust thanks to the different drag coefficients that a slender body experiences in the direction tangential and perpendicular to its motion in a viscous fluid [11, 227] (see Box III). A filamentous shape has been observed to be a viable strategy to avoid predation by protists by simply being too long to be ingested easily [319] (this approach is however futile at a certain size difference between devourer and the devoured, as is evident in the popularity of spaghetti as a human food source). The increased surface-to-volume ratio of a filamentous shape allows for easier uptake of nutrients [320], but also allows stronger attachment to substrates [321] in bacteria, because of a larger contact patch, compared to a spherical one.
@@
A filament’s most crucial mechanical property is its bending stiffness. It is known that one can draw conclusions about the forces at play from purely geometrical observations (see Box II). Techniques to measure the bending stiffness $\beta_{\mathrm{f}}$ and Young’s modulus $E$ of gliding filaments with the help of microfluidics [87] and micropipette techniques [89] have been developed (see Table 1).
@@
Filamentous cyanobacteria often exist in a surface-bound but motile state, called gliding [322, 323, 324]. The underlying mechanism that allows the generation of propulsive forces in this state is still a matter of debate [325, 326, 327, 328, 329], as no cyanobacteria species are known to possess flagella [323]. The main mechanisms that are debated are: type IV pili extension and retraction [330, 323], fibrils [329, 326], ejection of a jet of extracellular slime from junctional pores [331, 332, 333, 327] (at least for some species), and acoustic streaming [311, 328, 334]. By observing the self-buckling of a filament, the forces at play in its propulsion can be inferred via the bending stiffness [89]. Such observations combined with the independence of propulsion speed on filament length imply that all cells within a cyanobacteria filament contribute roughly equally to its propulsion.
@@
Cyanobacteria locomotion exhibits another interesting feature: path-tracking motion, where the body follows the path laid by the forward end of the filament [309, 310]. This type of locomotion in worms is also known as metameric locomotion [335]. Physically, information about a filament’s orientation persists over the time it takes a filament to traverse its own length [335]. Although their motility is induced by polar forces, filaments themselves are seemingly nematic, with no head or tail end. In fact, cyanobacteria have been observed to reverse their direction of motion frequently, as a strategy to navigate towards more favourable conditions, and even to form communities. [323, 336, 337, 338]. Given the delocalised nature of the filaments propulsive mechanism, a high degree of intercellular coordination is required to reverse the direction of motion, with the mechanisms of communication between cells not yet understood. Confined active filaments with low bending rigidity have been shown to curl up into tightly wound spirals (see Fig. 4(c)), effectively trapping themselves, greatly impeding their mobility [339, 340, 205, 341]. Utilising reversals is a strategy to escape this trapped state [307] as well as other forms of geometric confinement [342].
@@
Having a high aspect-ratio allows microbes to effectively sample changing conditions within space. While organisms with short aspect-ratio bodies commonly have to move around to find favourable conditions, essentially sampling the environment at random to inform their movement through e.g. chemo- or phototaxis, elongated filaments can sample gradients along their body and react before they have wholly immersed themselves in an unfavourable environment. Cyanobacteria, for instance, have been observed to react to light shining only onto parts of their body [343, 344], showing scotophobic responses, well before the whole filament loses access to light by reversing its gliding direction [214].
@@
Having an elongated body may even allow organisms to be in two places with different conditions at once. Cable bacteria are filamentous, multicellular bacteria belonging to the family Desulfobulbaceae. They can conduct electrons over centimetre scales [345, 346]. Through their filamentous shape they can couple sulfide oxidation in deeper, anoxic marine sediment with oxygen reduction in surface layers [347, 348, 349]. Electrical activity is also involved in a microbial feat of engineering: some electroactive bacteria produce nanowires to transfer electrons to extracellular electron acceptors [350]. These are, e.g., the Geobacter sulfurreducens or the Shewanella oneidensis MR-1 bacteria [351, 352, 353, 354].
@@
Active filaments can explore spaces inaccessible to plumper organisms with the same volume. Semiflexible active filaments have been shown to effectively explore complex and crowded environments [210, 209, 208, 355, 211]. The ability to reverse has been shown to further enhance the mobility of active filaments in porous environments [342]. Many soil dwelling microbes are filamentous in shape, in addition to their ability to explore a porous environment it has been speculated that their shape may enable them to fix their position in space, by wrapping around soil particles [320], allowing them to resist being washed away by percolating water. Entangled filamentous microbes have been observed to stabilize their habitats and protect them against erosion [356, 357]. Furthermore, nonmotile cells exploring a polymeric fluid can form long cables as they proliferate [358].
@@
Their filamentous nature allows filamentous organisms such as cyanobacteria to organise into structures much larger than an individual filament [309, 310, 359, 360]. Their filamentous shape allows them to glide over one another without losing contact with the substrate. The dynamics of gliding filaments result in the formation of a network-like reticulate pattern of filament bundles [309, 310, 360] (see Fig. 4(b)). Under different conditions, filaments have been observed to form aster-like aggregates [338]. These relatively simple, mostly two-dimensional structures may serve as a starting point for filaments to organize themselves into even more intricate three-dimensional structures, ranging from complex patterned biomats [359, 361] to the filaments weaving themselves into ropes [308] (see Fig. 4(d)), bridging large distances. The physical mechanisms behind this level of self-organisation and how the mechanical properties of these structures compare to that of a single filaments are yet to be understood, but can shed light on biofilm formation [362, 363]. It is plausible to presume that such complex assemblies could exhibit unusual rheological properties, just like those observed in cytoskeletal networks (see Sec. II).
@@ V Undulatory locomotion at the macroscopic scale
At the organismal scale, animals with a slender and deformable body shape are also ‘active filaments’, leading to the natural question of whether the analysis and framework presented in the previous sections could still apply. Diverse animals locomote by propagating undulatory waves. These body deformations are used to propel the organisms forward, although the actuation mechanisms may be very different across species and spatial scales (Figure 5). Well-studied examples include crawling worms [364, 365], swimming eels [366], and gliding snakes [367]. The study of both limbed and limbless movement has paved the way for understanding the origins and principles of neuro-mechanical control in diverse organisms, particularly the interactions between neural patterning and muscular activity [368]. Undulatory or sinusoidal locomotion strategies evolved across different phyla, often according to a basic neural circuit design involving mutually inhibitory halves with the left and right sides deforming alternately [369]. These insights into the neural basis of behaviour have led to the development of novel bio-inspired mechanisms for actuating synthetic systems including robots [370].
@@
Since propagating wave-like deformations are inherently non-reciprocal, they form the basis of diverse motility strategies through different media. For instance an undulating wave propagating head to tail down a sperm flagellum would break time reversibility since the filament will adopt a different sequence of shapes when time is reversed. However, non-reciprocity is a necessary but not sufficient condition [11] for net propulsion, as the latter also requires additional constraints on the number and placement of flagellar waves. Despite the large difference in scale compared to microscopic organisms such as flagellates and cyanobacteria (see Sec. III and IV), the undulatory locomotion of macroscopic animals such as nematodes and snakes can still be considered to operate in an overdamped regime where inertia is negligible, and where friction (instead of fluid viscosity) dominates. This is particularly true of locomotion in sand or granular media, where a resistive force theory based on drag anisotropy (see Box III) has been derived empirically, inspired by the movement of slender bodies through viscous fluids [371, 372].
@@
Figure 5: Undulatory locomotion of active filaments across scales in meso and macroscopic organisms. (a) C. elegans crawling movement on agar [373]. (b) Sidewinding locomotion of a snake, and (c) the timetrace of tracked points along the body of the snake as it moves on sand [374]. (d) Vinegar eels self-organise into metachronal waves (image courtesy of Alice Quillen [375]). (e) Topological entanglement in worm blobs [22]. (f) Bio-inspired snake-like robots with central-pattern-generator-like (CPG) control modules (see for example [376]).
@@
At the individual organism scale, distinct undulatory behaviours can arise from the same underlying structure and neural circuitry. A given body shape can be decomposed into a small number of principal components or oscillation modes, reflecting the intrinsically low-dimensional nature of the posture space, and allowing diverse undulatory shape modes to be reconstructed with only a small number of parameters [377]. In the example of C. elegans crawling, its motion in 2D is stereotyped and well-captured by just six basis modes, or eigenworms [365, 364]. Frequency and amplitude thus uniquely define a gait, which in turn determines the distance traversed by the organism in one cycle. Notions such as geometric phase can help understand and predict the mapping between these gaits and eventual displacement in the lab frame [378]. These dynamics are not totally unconstrained, however, undulatory locomotion across swimming sperm, nematodes, and even fish in overdamped environments may be subject to a universal dispersion relationship [379, 380]. Thus, active filaments can achieve highly reconfigurable dynamics with minimal neuromechanical coupling.
@@
Collectively, filamentous organisms also engage in novel behaviours that may not be accessible by individuals. At sufficiently high densities, these filaments self-assemble to form adaptive active materials with both solid-like and liquid-like properties, that are capable of emergent functions such as directed locomotion or smart navigation. Macroscopic ‘blobs’ of the 1 cm-long California blackworm Lumbriculus variegatus undergo dynamic collective movements to avoid desiccation or thermal stress without the need for any centralised control [381]. The entanglement, or ‘braiding’, process enables fast and reversible tangling and untangling of worm collectives. Intriguingly, L. variegatus blobs can take minutes to assemble but only milliseconds to disentangle. A recent study proposed an explanation for this phenomenon by experimentally imaging the 3D trajectories of the worms and mapping the dynamics onto a 3D filament model of Kirchhoff rods [22]. The unique, activity-dependent rheology of entangled blobs of Tubifex tubifex worms was also studied and shown to be shear-thinning, but with a different scaling compared to polymer melts [21]. These properties can even be harnessed to enable activity-dependent separation (chromatography) of worm mixtures [382]. For example, in T. tubifex the diffusivity is larger at low activity (i.e., low temperature) than at higher activity [208]. In another free-swimming nematode Turbatrix aceti (vingar eels), concentrated suspensions of these organisms display large-scale metachronal waves surprisingly similar to those found in ensembles of cilia and flagella [375, 383]. Understanding how these deformable filamentous bodies actively control the topology and rheology of the collective could inspire novel designs of functionally tunable synthetic materials [384].
@@
More generally, undulatory or anguilliform locomotion in elongated bodies allows for direct interactions between an organism and the surrounding medium, which can yield novel principles of biomimetic/bioinspired locomotion reliant on embodied design with few mechanical components. Self-propulsion can also be easily adapted for movement through different environments, for example for snakes the same basic locomotion mechanism can be equally effective for locomotion on land as well as in water [385, 386]. In more extreme cases, ‘flying’ snakes from the genus Chrysopelea use body undulations to generate lift to glide stably over several metres through the air [387].
@@
From a control theoretic perspective, undulatory locomotion is universal perhaps because it reduces the need for more complex neural control mechanisms and can be achieved simply by coordinating alternating waves of (usually muscle) contraction and relaxation. In some cases, oscillations emerge spontaneously as a result of instabilities, for example, dynamics reminiscent of ciliary beating have been observed in macroscopic chains of self-propelled microrobots, with the transition mapped to a supercritical Hopf bifurcation [388]. Indeed substrate-attached, microscopic elastic filaments forced at one end also exhibit oscillations in different regimes [389, 390, 391]. The propagation of waves thus provides one of the simplest possible appendage or body actuation mechanisms, and naturally breaks time-reversal symmetry (particularly relevant for inertia-free swimming). This results in highly adaptable and flexible control of movement where both the speed and direction of travel can be readily changed in response to stimuli.